if x=2+2^(2/3)+2^(1/3)
yhen x^3-6x^2+6x= ?
Please show the steps.

Grade:11

1 Answers

sanatan sharma
13 years ago

            x = 2 + 22/3 + 21/3

            x - 2 = 22/3 + 21/3

           cubing both the sides 

           (x - 2)3 = ( 22/3 + 21/3 )3

            x3 - 8 - 6x( x-2) = 4 + 2 + 6( x-2 )

            x3 - 8 -6x2 + 12x = 6 + 6x -12

             x3 -6x2 + 6x = 2 

          please approve the answer.